# The moduli spaces of left-invariant statistical structures on Lie groups

### v1

@@ 4.1. The moduli space of left-invariant Riemannian metrics
As is well-known, the Lie group $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ acts smoothly on the space of left-invariant Riemannian metrics $\mathop{\widetilde{\mathfrak{M}}}(G)$ (see Section  3.2 for more details on $\mathop{\widetilde{\mathfrak{M}}}(G)$) by
(4.1) $(r,\varphi).g\coloneqq r\cdot(\varphi^{-1})^{\ast}g=r\cdot g(\varphi^{-1}_{\ast}\ \cdot\ ,\varphi^{-1}_{\ast}\ \cdot\ )\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(G),\ g\in\mathop{\widetilde{\mathfrak{M}}}(G)).$
@@
(4.2) $(r,\varphi).\langle,\rangle\coloneqq r\cdot\langle\varphi^{-1}\ \cdot\ ,\varphi^{-1}\ \cdot\ \rangle\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(\mathfrak{g}),\ \langle,\rangle\in\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})).$
@@ 4.2. The moduli space of left-invariant statistical structures
(4.4) $(r,\varphi).(g,\nabla)\coloneqq(r\cdot(\varphi^{-1})^{\ast}g,(\varphi^{-1})^{\ast}\nabla)\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(G),\ (g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)).$
@@
Definition 4.2.1.
Let $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ be both left-invariant statistical structures on $G$. We say that a scaling statistical isomorphism $(f,r)$ from $(G,g,\nabla)$ to $(G,g^{\prime},\nabla^{\prime})$ (see Definition 2.2.1 for the notation) is compatible with the group structure if there exists $\varphi\in\mathrm{Aut}(G)$ such that $f(gx)=\varphi(g)f(x)$ for all $g,x\in G$.
@@
The proof of Proposition 4.2.2 is postponed to Appendix B (as Theorem B.0.3).
As a corollary to Proposition 4.2.2, one sees that the group action (4.4) preserves properties such as conjugate symmetry, constant curvature, dually flatness, and CHC for left-invariant statistical structures on $G$ (see Section 2.3 for the definitions of these classes).
Definition 4.2.3.
@@
It should be remarked that the moduli spaces, equipped with the quotient topology, might not be Hausdorff in general. In fact, the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$ is not Hausdorff if $G=\mathbb{R}^{n}$ (see Remark 5.1.6).
@@ 4.3. For the cases where $\mathfrak{PM}(G)\cong\{\ast\}$
By the general theory of the topology of equivariant fiber bundles on homogeneous spaces (cf. Theorem A.0.1 in Appendix A), we obtain the following:
Proposition 4.3.1.
@@
    is continuous, open and surjective. Here, $\nabla^{(g,C)}$ is the left-invariant statistical connection corresponding to $(g,C)$ in the sense of Proposition 3.2.1.
@@
(4.6) $(r,\varphi).C\coloneqq r\cdot C(\varphi^{-1}\ \cdot\ ,\varphi^{-1}\ \cdot\ ,\varphi^{-1}\ \cdot\ )\quad((r,\varphi)\in(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g},\ C\in S^{3}(\mathfrak{g}^{\ast})),$
@@
Proposition 4.3.2 ([37], Lemma 4.2).
@@ 5. The case of $\mathbb{R}^{n}$
In this section, we consider the commutative Lie group $\mathbb{R}^{n}$ with $n\geq 1$. Then $\mathfrak{PM}(\mathbb{R}^{n})\cong\{\ast\}$ (see Proposition 4.1.2).
@@ 5.1. Main theorem for $\mathbb{R}^{n}$
As a corollary to Theorem 5.1.1 (2), we have the following proposition:
Proposition 5.1.2.
@@
The Hessian structure $(g,\nabla^{g})$ is known to be of dually flat of CHC 0 on $\mathbb{R}^{n}$ (see [34]). Conversely, we suppose that $(g,\nabla)$ is a left-invariant dually flat structure on $\mathbb{R}^{n}$ of CHC, and let us prove that $\nabla=\nabla^{g}$. By Proposition 2.3.2 and $R^{\nabla}=R^{\nabla^{g}}=0$, we have $K_{X},K_{Y}]=0$ for any $X,Y\in\mathfrak{g}_{\mathbb{R}^{n}}$. Here, $K$ is the difference tensor of $(g,\nabla)$. We also note that $(\mathbb{R}^{n},g)$ has the constant sectional curvature 0. Thus by [34, Corollary 2] and [10, Proposition 3.1], for each $X,Y,Z\in\mathfrak{g}_{\mathbb{R}^{n}}$, the equality below holds:
@@
Take the orthonormal basis $e_{1},\dots,e_{n}$ of $\mathfrak{g}_{\mathbb{R}^{n}}$ as in Theorem 5.1.1 (2) and write $C=\sum_{i}\lambda_{i}x_{i}^{3}$. Note that $K(e_{i},e_{i})=\lambda_{i}e_{i}$ for each $i$. Therefore, we have
@@
$\Omega_{c}:=\{(x_{1},\dots,x_{n})\mid x_{i}>0\text{ for each }i\leq p\},$
@@
By Proposition 4.1.2, and Theorem 5.1.1, there exists an orthogonal basis $\mathcal{B}=\{e_{1},\dots,e_{n}\}$ on $(\mathfrak{g}_{\mathbb{R}^{n}},\langle,\rangle)$ and $c=(c_{1},\dots,c_{n})\in\{\lambda\in\mathbb{R}^{n}\mid\lambda_{1}\geq\dots\lambda_{n}\geq 0\}$ such that by putting $C=\sum_{i}c_{i}x_{i}^{3}$, the dually flat statistical manifold $(\mathbb{R}^{n},g,\nabla)$ is statistically isomorphic to $(\mathbb{R}^{n},g^{E},C)$ up to scaling. Then, it can be directly verified that the diffeomorphism $\phi_{c}$ gives a scaling statistical isomorphism from the dually flat statistical manifold $(\Omega_{c},g_{c},\nabla_{c})$ onto $(\mathbb{R}^{n},g^{E},C)$. ∎
Remark 5.1.5.
It should be noted that, in the context of Hessian geometry, the claim of Corollary 5.1.4 can essentially be found in [33, 35], where it appears as the classification of homogeneous Hessian domains associated with commutative normal Hessian algebras.
Remark 5.1.6.
@@ 5.2. Left-invariant conjugate-symmetric and dually flat statistical structures on $\mathbb{R}^{n}$
In this subsection, we give a proof of Theorem 5.1.1.
Proof of (1) in Theorem 5.1.1.
By Proposition 4.1.2, our goal is to show that $S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast},\langle,\rangle)=S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$. Fix any $C\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$. Since $\nabla^{g^{E}}_{Y}X=0$ for any $X,Y\in\mathfrak{g}_{\mathbb{R}^{n}}^{\ast}$, we have $\nabla^{g^{E}}C=0$. This proves that $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast},\langle,\rangle)$ by Proposition 2.3.1. ∎
Proof of (2) in Theorem 5.1.1.
Let $C\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$. Throughout this proof, the difference tensor of the left-invariant statistical structure $(\langle,\rangle,C)$ is denoted by $K$, and the affine connection by $\nabla$ in the sense of Proposition 3.2.1. Moreover, we denote the $(1,3)$-curvature tensor of $\nabla$ by $R$, and the $(1,3)$-curvature tensor of Levi-Civita connection of $g^{E}$ by $R_{g^{E}}\equiv 0$.
@@
Assume (ii). Then by Proposition 3.2.1, the difference tensor $K$ satisfies
@@
where $\delta_{ij}$ denotes the Kronecker delta. We aim to show that $R$ coincides with 0. By combining $R_{g^{E}}\equiv 0$, Theorem 5.1.1 (1) with the equality Proposition 2.3.2, we see the following.
@@
By Equation (5.2) and Proposition 2.2.4, we see that
@@
Assume (i). Recall that any dually flat structure is conjugate symmetric. Thus, by Proposition 2.3.2 and $R_{g^{E}}\equiv 0$, we have the equality
@@
holds for all $u$, $v\in\mathfrak{g}_{\mathbb{R}^{n}}$. By Proposition 2.2.4, there exists an orthonormal basis $e_{1},\dots,e_{n}$ of $(\mathfrak{g}_{\mathbb{R}^{n}},\langle,\rangle)$ and $c_{1},\dots c_{n}\in\mathbb{R}$ such that
@@
follows from Proposition 4.3.1, Proposition 4.3.2 (1) and Theorem 5.1.1 (1). The claim
@@
is obtained by applying Proposition 4.3.1, Proposition 4.3.2 (1), Theorem 5.1.1 (2), and Theorem C.0.1 in Appendix C. ∎
@@ 6. The case of $G_{\mathbb{R}\mathrm{H}^{n}}$
In this section, we consider the connected and simply-connected Lie group $G_{\mathbb{R}\mathrm{H}^{n}}$ with $n\geq 2$ (see Section 4.1 for the definition of $G_{\mathbb{R}\mathrm{H}^{n}}$). Then $\mathfrak{PM}(G_{\mathbb{R}\mathrm{H}^{n}})\cong\{\ast\}$ (see Proposition 4.1.2).
@@ 6.1. Main theorem for $G_{\mathbb{R}\mathrm{H}^{n}}$
Remark 6.1.2.
The dually flat structure $(g,C^{1})$ can be found in [8, Example 2.3], as an example of Hessian structures of CHC $4$ on hyperbolic spaces. Notice that its dual structure $(\langle,\rangle,C^{-1})$ on $G_{\mathbb{R}\mathrm{H}^{n}}$ is dually flat but not CHC.
As supplementary remarks to Theorem 6.1.1, we will also discuss the following points in Section 6.3 (see Theorem 6.3.1):
Proposition 6.1.3.
@@
In this subsection, we give a proof of Theorem 6.1.1.
@@
Since $\mathfrak{PM}(G_{\mathbb{R}\mathrm{H}^{n}})\cong\{\ast\}$, by applying Proposition 4.3.1 and 4.3.2 (2), we have
@@
Therefore Theorem 6.1.1 follows from the three propositions below.
Proposition 6.2.1.
@@
First, we shall give a proof of Proposition 6.2.1:
Proof of Proposition 6.2.1.
@@
In order to prove Proposition 6.2.2, we prepare the following lemma:
Lemma 6.2.4.
@@
Proof of Lemma 6.2.4.
By definition, the structure constants of $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$ with respect to the canonical basis are given by $a_{1t}^{t}=1$, $a_{t1}^{t}=-1$ ($t\geq 2$), with all other $a_{ij}^{k}$ equal to zero. Then, by applying Proposition 3.3.1, the generalized Christoffel symbols can be computed as
@@
with all other $\Gamma_{ij}^{k}$ equal to zero. The claims of Lemma 6.2.4 can be obtained by applying Proposition 3.3.2. ∎
Let us give a proof of Proposition 6.2.2.
Proof of Proposition 6.2.2.
@@
and $C=\sum_{u,v,s}C_{uvs}(x_{uvs})$, where we put $x_{uvs}:=x_{u}\otimes x_{v}\otimes x_{s}\in(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})^{\otimes 3}$ and $x_{uvst}:=x_{u}\otimes x_{v}\otimes x_{s}\otimes x_{t}\in(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})^{\otimes 4}$. Then by Lemma 6.2.4,
@@
Thus $C=C_{111}(x_{111}+\frac{3}{2}\sum_{i\geq 2}x_{1}x_{i}^{2})=C^{\alpha}$ for $\alpha=C_{111}/4$. Conversely, by Lemma 6.2.4, one can see that
@@
Finally, we shall prove Proposition 6.2.3:
Proof of Proposition 6.2.3.
Fix $\alpha\in\mathbb{R}$, and put $C:=C^{\alpha}$. Recall that $R^{0}_{g}=\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}$. Thus by applying Proposition 2.3.2, our goal is to show that
@@
Therefore by Theorem 3.3.3,
$[K,K]$ $=(-\alpha^{2})\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}=(-\alpha^{2})R^{0}_{g}.$

### v3

@@ 4.1. The moduli space of left-invariant Riemannian metrics
As is well-known, the Lie group $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ acts smoothly on the space of left-invariant Riemannian metrics $\mathop{\widetilde{\mathfrak{M}}}(G)$ (see Section 3.2 for more details on $\mathop{\widetilde{\mathfrak{M}}}(G)$) by
(4.1) $(r,\varphi).g\coloneqq r\cdot(\varphi^{-1})^{\ast}g=r\cdot g(\varphi^{-1}_{\ast}~\cdot~,\varphi^{-1}_{\ast}~\cdot~)\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(G),~g\in\mathop{\widetilde{\mathfrak{M}}}(G)).$
@@
(4.2) $(r,\varphi).\langle,\rangle\coloneqq r\cdot\langle\varphi^{-1}~\cdot~,\varphi^{-1}~\cdot~\rangle\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(\mathfrak{g}),~\langle,\rangle\in\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})).$
@@ 4.2. The moduli space of left-invariant statistical structures
(4.4) $(r,\varphi).(g,\nabla)\coloneqq(r\cdot(\varphi^{-1})^{\ast}g,(\varphi^{-1})^{\ast}\nabla)\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(G),~(g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)).$
@@
Definition 4.3.
Let $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ be both left-invariant statistical structures on $G$. We say that a scaling statistical isomorphism $(f,r)$ from $(G,g,\nabla)$ to $(G,g^{\prime},\nabla^{\prime})$ (see Definition 2.2 for the notation) is compatible with the group structure if there exists $\varphi\in\mathrm{Aut}(G)$ such that $f(gx)=\varphi(g)f(x)$ for all $g,x\in G$.
@@
The proof of Proposition 4.4 is postponed to Appendix B (as Theorem B.3).
As a corollary to Proposition 4.4, one sees that the group action (4.4) preserves properties such as conjugate symmetry, constant curvature, dually flatness, and CHC for left-invariant statistical structures on $G$ (see Section 2.3 for the definitions of these classes).
Definition 4.5.
@@
It should be remarked that the moduli spaces, equipped with the quotient topology, might not be Hausdorff in general. In fact, the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$ is not Hausdorff if $G=\mathbb{R}^{n}$ (see Remark 5.6).
@@ 4.3. For the cases where $\mathfrak{PM}(G)\cong\{\ast\}$
By the general theory of the topology of equivariant fiber bundles on homogeneous spaces (cf. Theorem A.1 in Appendix A), we obtain the following:
Proposition 4.6.
@@
    is continuous, open and surjective. Here, $\nabla^{(g,C)}$ is the left-invariant statistical connection corresponding to $(g,C)$ in the sense of Proposition 3.3.
@@
(4.6) $(r,\varphi).C\coloneqq r\cdot C(\varphi^{-1}~\cdot~,\varphi^{-1}~\cdot~,\varphi^{-1}~\cdot~)\quad((r,\varphi)\in(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g},~C\in S^{3}(\mathfrak{g}^{\ast})),$
@@
Proposition 4.7 (cf. Lemma 4.2 in Taketomi (2022), arXiv:2210.01483).
@@ 5. The case of $\mathbb{R}^{n}$
In this section, we consider the commutative Lie group $\mathbb{R}^{n}$ with $n\geq 1$. Then $\mathfrak{PM}(\mathbb{R}^{n})\cong\{\ast\}$ (see Proposition 4.2).
@@ 5.1. Main theorem for $\mathbb{R}^{n}$
As a corollary to Theorem 5.1 (2), we have the following proposition:
Proposition 5.2.
@@
The Hessian structure $(g,\nabla^{g})$ is known to be of dually flat of CHC 0 on $\mathbb{R}^{n}$ (see [33]). Conversely, we suppose that $(g,\nabla)$ is a left-invariant dually flat structure on $\mathbb{R}^{n}$ of CHC, and let us prove that $\nabla=\nabla^{g}$. By Proposition 2.7 and $R^{\nabla}=R^{\nabla^{g}}=0$, we have $K_{X},K_{Y}]=0$ for any $X,Y\in\mathfrak{g}_{\mathbb{R}^{n}}$. Here, $K$ is the difference tensor of $(g,\nabla)$. We also note that $(\mathbb{R}^{n},g)$ has the constant sectional curvature 0. Thus by [33, Corollary 2] and [10, Proposition 3.1], for each $X,Y,Z\in\mathfrak{g}_{\mathbb{R}^{n}}$, the equality below holds:
@@
Take the orthonormal basis $e_{1},\dots,e_{n}$ of $\mathfrak{g}_{\mathbb{R}^{n}}$ as in Theorem 5.1 (2) and write $C=\sum_{i}\lambda_{i}x_{i}^{3}$. Note that $K(e_{i},e_{i})=\lambda_{i}e_{i}$ for each $i$. Therefore, we have
@@
$\Omega_{c}:=\left\{(x_{1},\dots,x_{n})~\middle|~x_{i}>0\text{ for each }i\leq p\right\},$
@@
By Proposition 4.2, and Theorem 5.1, there exists an orthogonal basis $\mathcal{B}=\{e_{1},\dots,e_{n}\}$ on $(\mathfrak{g}_{\mathbb{R}^{n}},\langle,\rangle)$ and $c=(c_{1},\dots,c_{n})\in\{\lambda\in\mathbb{R}^{n}\mid\lambda_{1}\geq\dots\lambda_{n}\geq 0\}$ such that by putting $C=\sum_{i}c_{i}x_{i}^{3}$, the dually flat statistical manifold $(\mathbb{R}^{n},g,\nabla)$ is statistically isomorphic to $(\mathbb{R}^{n},g^{E},C)$ up to scaling. Then, it can be directly verified that the diffeomorphism $\phi_{c}$ gives a scaling statistical isomorphism from the dually flat statistical manifold $(\Omega_{c},g_{c},\nabla_{c})$ onto $(\mathbb{R}^{n},g^{E},C)$. ∎
Remark 5.5.
It should be noted that, in the context of Hessian geometry, the claim of Corollary 5.4 can essentially be found in [32, 34], where it appears as the classification of homogeneous Hessian domains associated with commutative normal Hessian algebras.
Remark 5.6.
@@ 5.2. Left-invariant conjugate-symmetric and dually flat statistical structures on $\mathbb{R}^{n}$
In this subsection, we give a proof of Theorem 5.1.
Proof of (1) in Theorem 5.1.
By Proposition 4.2, our goal is to show that $S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast},\langle,\rangle)=S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$. Fix any $C\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$. Since $\nabla^{g^{E}}_{Y}X=0$ for any $X,Y\in\mathfrak{g}_{\mathbb{R}^{n}}^{\ast}$, we have $\nabla^{g^{E}}C=0$. This proves that $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast},\langle,\rangle)$ by Proposition 2.6. ∎
Proof of (2) in Theorem 5.1.
Let $C\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$. Throughout this proof, the difference tensor of the left-invariant statistical structure $(\langle,\rangle,C)$ is denoted by $K$, and the affine connection by $\nabla$ in the sense of Proposition 3.3. Moreover, we denote the $(1,3)$-curvature tensor of $\nabla$ by $R$, and the $(1,3)$-curvature tensor of Levi-Civita connection of $g^{E}$ by $R_{g^{E}}\equiv 0$.
@@
Assume (ii). Then by Proposition 3.3, the difference tensor $K$ satisfies
@@
where $\delta_{ij}$ denotes the Kronecker delta. We aim to show that $R$ coincides with 0. By combining $R_{g^{E}}\equiv 0$, Theorem 5.1 (1) with the equality Proposition 2.7, we see the following.
@@
By Equation (5.2) and Proposition 2.5, we see that
@@
Assume (i). Recall that any dually flat structure is conjugate symmetric. Thus, by Proposition 2.7 and $R_{g^{E}}\equiv 0$, we have the equality
@@
holds for all $u$, $v\in\mathfrak{g}_{\mathbb{R}^{n}}$. By Proposition 2.5, there exists an orthonormal basis $e_{1},\dots,e_{n}$ of $(\mathfrak{g}_{\mathbb{R}^{n}},\langle,\rangle)$ and $c_{1},\dots c_{n}\in\mathbb{R}$ such that
@@
follows from Proposition 4.6, Proposition 4.7 (1) and Theorem 5.1 (1). The claim
@@
is obtained by applying Proposition 4.6, Proposition 4.7 (1), Theorem 5.1 (2), and Theorem C.1 in Appendix C. ∎
@@ 6. The case of $G_{\mathbb{R}\mathrm{H}^{n}}$
In this section, we consider the connected and simply-connected Lie group $G_{\mathbb{R}\mathrm{H}^{n}}$ with $n\geq 2$ (see Section 4.1 for the definition of $G_{\mathbb{R}\mathrm{H}^{n}}$). Then $\mathfrak{PM}(G_{\mathbb{R}\mathrm{H}^{n}})\cong\{\ast\}$ (see Proposition 4.2).
@@ 6.1. Main theorem for $G_{\mathbb{R}\mathrm{H}^{n}}$
Remark 6.2.
The dually flat structure $(g,C^{-1})$ can be found in [8, Example 2.3], as an example of Hessian structures of CHC $4$ on hyperbolic spaces. Notice that its dual structure $(g,C^{1})$ on $G_{\mathbb{R}\mathrm{H}^{n}}$ is dually flat but not CHC.
As supplementary remarks to Theorem 6.1, we will also discuss the following points in Section 6.3 (see Theorem 6.8):
Proposition 6.3.
@@
In this subsection, we give a proof of Theorem 6.1.
@@
Since $\mathfrak{PM}(G_{\mathbb{R}\mathrm{H}^{n}})\cong\{\ast\}$, by applying Propositions 4.6 and 4.7 (2), we have
@@
Therefore Theorem 6.1 follows from the three propositions below.
Proposition 6.4.
@@
First, we shall give a proof of Proposition 6.4:
Proof of Proposition 6.4.
@@
In order to prove Proposition 6.5, we prepare the following lemma:
Lemma 6.7.
@@
Proof of Lemma 6.7.
By definition, the structure constants of $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$ with respect to the canonical basis are given by $a_{1t}^{t}=1$, $a_{t1}^{t}=-1$ ($t\geq 2$), with all other $a_{ij}^{k}$ equal to zero. Then, by applying Proposition 3.4, the generalized Christoffel symbols can be computed as
@@
with all other $\Gamma_{ij}^{k}$ equal to zero. The claims of Lemma 6.7 can be obtained by applying Proposition 3.5. ∎
Let us give a proof of Proposition 6.5.
Proof of Proposition 6.5.
@@
and $C=\sum_{u,v,s}C_{uvs}(x_{uvs})$, where we put $x_{uvs}:=x_{u}\otimes x_{v}\otimes x_{s}\in(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})^{\otimes 3}$ and $x_{uvst}:=x_{u}\otimes x_{v}\otimes x_{s}\otimes x_{t}\in(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})^{\otimes 4}$. Then by Lemma 6.7,
@@
Thus $C=C_{111}(x_{111}+\frac{3}{2}\sum_{i\geq 2}x_{1}x_{i}^{2})=C^{\alpha}$ for $\alpha=C_{111}/4$. Conversely, by Lemma 6.7, one can see that
@@
Finally, we shall prove Proposition 6.6:
Proof of Proposition 6.6.
Fix $\alpha\in\mathbb{R}$, and put $C:=C^{\alpha}$. Recall that $R^{0}_{g}=\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}$. Thus by applying Proposition 2.7, our goal is to show that
@@
Therefore by Theorem 3.6,
$K,K]=(-\alpha^{2})\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}=(-\alpha^{2})R^{0}_{g}.$